OpenStudy (anonymous):

How do you solve: What is the molarity of the mixture obtained from mixing 1.250 L of 0.040 M solution and 750 mL of 0.20 M? The answer is 0.100 M.

Vocaloid (vocaloid):

calculate the number of moles in each of the two solutions, then divide by the total number of liters moles = (volume)(molarity) number of moles in the first solution = (1.250)(0.040) = 0.05 moles number of moles in the second solution = (.750)(0.20) = 0.15 moles total number of liters = .750 + 1.250 = 2 L moles/liters = (0.05+0.15)/2L = 0.1 M